# PACS Collaboration

### v1

@@ Abstract
We performed a precise calculation of physical quantities related to the axial structure of the nucleon using 2+1 flavor lattice QCD gauge configuration (PACS10 configuration) generated at the physical point with lattice volume larger than $(10\;{\mathrm{fm}})^{4}$ by the PACS Collaboration. The nucleon matrix element of the axial-vector current has two types of the nucleon form factors, the axial-vector ($F_{A}$) form factor and the induced pseudoscalar ($F_{P}$) form factor. Recently lattice QCD simulations have succeeded in reproducing the experimental value of the axial-vector coupling, $g_{A}$, determined from $F_{A}(q^{2})$ at zero momentum transfer $q^{2}=0$, at a percent level of statistical accuracy. However, the $F_{P}$ form factor so far has not reproduced the experimental values well due to strong $\pi N$ excited-state contamination. Therefore, we proposed a simple subtraction method for removing the so-called leading $\pi N$-state contribution, and succeeded in reproducing the values obtained by two experiments of muon capture on the proton and pion electro-production for $F_{P}(q^{2})$. The novel approach can also be applied to the nucleon pseudoscalar matrix element to determine the pseudoscalar ($G_{P}$) form factor with the help of the axial Wald-Takahashi identity. The resulting form factors, $F_{P}(q^{2})$ and $G_{P}(q^{2})$, are in good agreement with the prediction of the pion-pole dominance model. In the new analysis, the induced pseudoscalar coupling $g_{P}^{\ast}$ and the pion-nucleon coupling $g_{\pi NN}$ can be evaluated with a few percent accuracy including systematic uncertainties using existing data calculated at two lattice spacings.
@@ I Introduction
The axial structure of the nucleon plays an important role to understand the weak interaction between nucleons. Experimentally, weak processes meditated by the weak charged current like neutron beta decay $n\rightarrow p+e^{-}+\bar{\nu}_{e}$, muon capture on the proton $\mu^{-}+p\rightarrow\nu_{\mu}+n$, or quasi elastic neutrino scattering $\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+n$ are mainly exploited for studying the weak nucleon form factors. In recent years, the physics of neutrino oscillations, which inevitably involves the effects of nucleons and nuclear structure, has required accurate knowledge of the axial structure of the nucleon Kronfeld:2019nfb ; Meyer:2022mix ; Tomalak:2023pdi ; Petti:2023abz . This is because the vector part of weak processes can be constrained by knowledge of the isovector combination between the electromagnetic form factors of the proton and the neutron under the assumption of the conserved-vector-current (CVC) hypothesis using isospin relations Thomas:2001kw , while the axial-vector part described by the axial-vector ($F_{A}$) and induced pseudoscalar ($F_{P}$) form factors is less known except for the axial-vector coupling $g_{A}=F_{A}(q^{2}=0)$ that is most accurately measured by neutron beta decay.
From a theoretical perspective, the axial structure of the nucleon is highly connected with the physics of chiral symmetry and its spontaneous breaking, which ensures the presence of pseudo Nambu-Goldstone particles such as the pion. This is empirically known as the partially conserved axial-vector current (PCAC) hypothesis, where the divergence of the axial-vector current is proportional to the pion field. Applying this idea to the isovector axial-vector matrix element of the nucleon given by
@@
which is satisfied among the three nucleon form factors including the pesudoscalar ($G_{P}$) form factor defined in the isovector pseudoscalar matrix element of the nucleon as
@@
We have extensively studied the axial structure of the nucleon in lattice QCD at the physical point using two sets of the PACS10 gauge configurations  [^1] In the PACS10 project, the PACS Collaboration have generated three sets of the PACS10 gauge configurations at the physical point with lattice volume larger than $(10\;{\rm fm})^{4}$ and three different lattice spacings Yamazaki:2024otm . Our study using the third PACS10 gauge configurations is currently underway Tsuji:2024scy . in our previous works Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut ; Tsuji:2022ric , which perform lattice QCD simulations as listed in Table 1. In our strategy, the smearing parameters of the nucleon interpolation operator were highly optimized to eliminate as much as possible the contribution of excited states in the nucleon two-point function. This strategy was quite successful in calculations of the $F_{A}$ form factor, while the $F_{P}$ and $G_{P}$ form factors remained strongly affected by residual contamination of $\pi N$-state contribution. As results, both the $F_{P}$ and $G_{P}$ form factors are significantly underestimated in the low-$q^{2}$ region compared to the PPD model and do not satisfy the generalized GT relation Shintani:2018ozy ; Tsuji:2023llh .
In this work, we propose a simple method to remove the $\pi N$-state contamination from the $F_{P}$ and $G_{P}$ form factors given by the standard ratio method in lattice QCD Sasaki:2025qro . The application of the novel method will enable the precise determination of the induced pseudoscalar coupling ($g_{P}^{\ast}$) and the pion-nucleon coupling ($g^{\pi NN}$) from existing data calculated at two lattice spacings, since these quantities are defined using the $F_{P}$ form factor Thomas:2001kw as
@@
$g_{\pi NN}$ $=\lim_{q^{2}\rightarrow\infty}(q^{2}+M_{\pi}^{2})\frac{F_{P}(q^{2})}{2F_{\pi}}$
@@
This paper is organized as follows. In Sec. II, after a brief introduction to the standard ratio method for calculating the nucleon form factors in lattice QCD simulations, we describe our proposal to remove $\pi N$-state contamination from two specific nucleon form factors $F_{P}(q^{2})$ and $G_{P}(q^{2})$ (hereafter referred to as the leading $\pi N$ subtraction method). Section III presents the numerical results of the $F_{P}$ and $G_{P}$ form factors obtained by the leading $\pi N$ subtraction method with the data sets generated in Refs. Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut . We then determine the values of $g_{P}^{\ast}$ and $g_{\pi NN}$. Finally, we close with a summary in Sec. IV.
In this paper, the nucleon matrix elements are given in the Euclidean metric convention. $\gamma_{5}$ is defined by $\gamma_{5}\equiv\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}=-\gamma_{5}^{M}$, which has the opposite sign relative to that in the Minkowski convention ($\vec{\gamma}^{M}=i\vec{\gamma}$ and $\gamma^{M}_{0}=\gamma_{4}$) adopted in the particle data group. The sign of all the form factors is chosen to be positive. The Euclidean four momentum squared $q^{2}$ corresponds to the spacelike momentum squared as $q_{M}^{2}=-q^{2}<0$ in Minkowski space.
Table 1: Details of gauge ensembles: the $\beta$ value, lattice volume ($L^{3}\times T$), spatial extent ($La$), the simulated pion mass ($M_{\pi}$) and the number of configurations ($N_{\mathrm{conf}}$) to be analyzed for studying the isovector nucleon form factors in our previous works Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut ; Tsuji:2022ric and also this study.
$\beta$ Label $L^{3}\times T$ $La$ [fm] $M_{\pi}$ [GeV] $N_{\mathrm{conf}}$
1.82 PACS10/L128 $128^{3}\times 128$ 10.8 0.135 20
PACS5/L64 $64^{3}\times 64$ 5.9 0.138 100
2.00 PACS10/L160 $160^{3}\times 160$ 10.1 0.138 19
@@ II.1 Standard ratio method
where the local current $\widetilde{J}$ 222 Hereafter, the currents and the form factors with and without tilde indicate bare and renormalized ones. located at the time slice $t$ denotes the axial-vector current ($J=A_{\alpha}$) or the pseudoscalar current ($J=P$), carrying the momentum transfer $\bm{q}=\bm{p}-\bm{p}^{\prime}$. The projection operator ${\cal P}_{+}=\frac{1+\gamma_{4}}{2}$ appearing in Eq. (7) can eliminate contributions from the opposite-parity state for the case of $|\bm{p}|=0$ Sasaki:2001nf ; Sasaki:2005ug , while the projection operator ${\cal P}^{5z}={\cal P}_{+}\gamma_{5}\gamma_{3}$ appearing in Eq. (8) means that the $z$ direction is chosen as the polarization direction.
In a conventional way to extract the nucleon form factors, the following ratios are constructed by appropriate combinations of two-point and three-point functions with the source-sink separation ($t_{\mathrm{sep}}\equiv t_{\mathrm{sink}}-t_{\mathrm{src}}$agler:2003jd ; Gockeler:2003ay as
@@
Figure 1: Using the standard ratio method, the values of $F_{A}$ (upper-left panel), $2M_{N}F_{P}$ (upper-right panel) and $\widetilde{G}_{P}$ (lower panel) calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven non-zero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$.
The ratio $\mathcal{R}_{J}^{5z}(t;\bm{q})$ gives the following asymptotic values including the respective bare form factors in the asymptotic region Sasaki:2007gw :
@@
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh the correlation functions involving the $A_{4}$ current are not taken into account for the calculation of the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ form factors. Instead of using Eq. (13), kinematically different types of the three-point functions (12) are used to extract the two independent form factors $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$, since the $z$ direction is chosen as the polarization direction by the projection operator ${\cal P}^{5z}$ Sasaki:2007gw . Indeed, the longitudinal momentum ($q_{3}$) dependence explicitly appears in Eq. (12). Therefore, in the transverse case ($i=1$ or 2), ${\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ only contains the $\widetilde{F}_{P}(q^{2})$ form factor, while ${\cal R}^{5z}_{A_{3}}(t,{\bm{q}})$ contains both $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ contributions as shown in Eq. (12). Even in the longitudinal case ($i=3$), the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ can be separately determined due to the fact that there are two types of kinematics: $q_{3}=0$ or $q_{3}\neq 0$. For the former case, ${\cal R}^{5z}_{A_{3}}(t,{\bm{q}})$ only contains the $\widetilde{F}_{A}(q^{2})$. Therefore, the subtracted correlator ratio ¯$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ that only includes the $\widetilde{F}_{P}(q^{2})$ contribution can be defined as
¯$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ $\equiv{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})-\delta_{i3}{\cal R}^{5z}_{A_{3}}(t,{\bm{q}}_{0})$ (15)
@@
Finally, we recall that the quark local currents on the lattice receive finite renormalizations relative to their continuum counterparts in general. The renormalized values of the form factors require the renormalization factors $Z_{O}\ (O=A,P)$:
@@
where the renormalization factors are defined through the renormalization of the quark local currents, $A_{\alpha}=Z_{A}\widetilde{A}_{\alpha}$ and $P=Z_{P}\widetilde{P}$. The renormalization factor $Z_{A}$ is scale independent, while $Z_{P}$ depends on the renormalization scale. To make a comparison with the experimental values, two form factors, $F_{A}$ and $F_{P}$, will be properly renormalized with $Z_{A}$, which is determined by the Schrödinger functional method, while the pseudoscalar form factor presented in this study is only bare quantity as indicated by $\widetilde{G}_{P}$.
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , we adopted the strategy that the smearing parameters of the nucleon interpolating operator are highly optimized to eliminate as much as possible the contribution of excited states in the nucleon two-point function. As shown in Fig. 1, the values of $F_{A}$ (upper-left panel), $2M_{N}F_{P}$ (upper-right panel) and $\widetilde{G}_{P}$ (lower panel) computed on the $160^{4}$ lattice (PACS10/L160) with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven non-zero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. Although $F_{A}(q^{2})$ shows a clear plateau and no $t_{\mathrm{sep}}$ dependence, suggesting that the ground-state dominance is manifest thanks to the elaborate tuning of the sink and source smearing functions, the excited-state contributions are not completely eliminated in $F_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$.

### v2

@@ Abstract
We perform a precise calculation of physical quantities related to the axial structure of the nucleon using 2+1 flavor lattice QCD gauge configuration (PACS10 configuration) generated at the physical point with lattice volume larger than $(10\;{\mathrm{fm}})^{4}$ by the PACS Collaboration. The nucleon matrix element of the axial-vector current has two types of the nucleon form factors, the axial-vector ($F_{A}$) form factor and the induced pseudoscalar ($F_{P}$) form factor. Recently lattice QCD simulations have succeeded in reproducing the experimental value of the axial-vector coupling, $g_{A}$, determined from $F_{A}(q^{2})$ at zero momentum transfer $q^{2}=0$, at a percent level of statistical accuracy. However, the $F_{P}$ form factor so far has not reproduced the experimental values well due to strong $\pi N$ excited-state contamination. Therefore, we propose a simple subtraction method for removing the so-called leading $\pi N$-state contribution. This method succeeds in reproducing the values obtained by two experiments of muon capture on the proton and pion electro-production for $F_{P}(q^{2})$. The novel approach can also be applied to determine the pseudoscalar ($G_{P}$) form factor from the nucleon pseudoscalar matrix element with the help of the axial Ward-Takahashi identity. The resulting form factors, $F_{P}(q^{2})$ and $G_{P}(q^{2})$, are in good agreement with the prediction of the pion-pole dominance model. In the new analysis, the induced pseudoscalar coupling $g_{P}^{\ast}$ and the pion-nucleon coupling $g_{\pi NN}$ can be evaluated with a few percent accuracy including systematic uncertainties using existing data calculated at two lattice spacings.
@@ I Introduction
The axial structure of the nucleon plays an important role to understand the weak interaction between nucleons. Experimentally, weak processes mediated by the weak charged current like neutron beta decay $n\rightarrow p+e^{-}+\bar{\nu}_{e}$, muon capture on the proton $\mu^{-}+p\rightarrow\nu_{\mu}+n$, or quasi-elastic neutrino scattering $\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+n$ are mainly exploited for studying the weak nucleon form factors. In recent years, the physics of neutrino oscillations, which inevitably involves the effects of nucleons and nuclear structure, has required accurate knowledge of the axial structure of the nucleon Kronfeld:2019nfb ; Meyer:2022mix ; Tomalak:2023pdi ; Petti:2023abz . This is because the vector part of weak processes can be constrained by knowledge of the isovector combination between the electromagnetic form factors of the proton and the neutron under the assumption of the conserved-vector-current (CVC) hypothesis using isospin relations Thomas:2001kw , while the axial-vector part described by the axial-vector ($F_{A}$) and induced pseudoscalar ($F_{P}$) form factors is less known except for the axial-vector coupling $g_{A}=F_{A}(q^{2}=0)$ that is most accurately measured by neutron beta decay.
From a theoretical perspective, the axial structure of the nucleon is highly connected with the physics of chiral symmetry and its spontaneous breaking, which ensures the presence of pseudo-Nambu-Goldstone particles such as the pion. This is empirically known as the partially conserved axial-vector current (PCAC) hypothesis, where the divergence of the axial-vector current is proportional to the pion field. Applying this idea to the isovector axial-vector matrix element of the nucleon given by
@@
which is satisfied among the three nucleon form factors including the pseudoscalar ($G_{P}$) form factor defined in the isovector pseudoscalar matrix element of the nucleon as
@@
We have extensively studied the axial structure of the nucleon in lattice QCD at the physical point using two sets of the PACS10 gauge configurations  [^1] In the PACS10 project, the PACS Collaboration have generated three sets of the PACS10 gauge configurations at the physical point with lattice volume larger than $(10\;{\rm fm})^{4}$ and three different lattice spacings Yamazaki:2024otm . Our study using the third PACS10 gauge configurations is underway Tsuji:2024scy . in our previous works Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut ; Tsuji:2022ric , which perform lattice QCD simulations as listed in Table 1. In our strategy, the smearing parameters of the nucleon interpolation operator were highly optimized to eliminate as much as possible the contribution of excited states in the nucleon two-point function. This strategy was quite successful in calculations of the $F_{A}$ form factor, while the $F_{P}$ and $G_{P}$ form factors remained strongly affected by residual contamination of $\pi N$-state contribution. As a result, both the $F_{P}$ and $G_{P}$ form factors are significantly underestimated in the low-$q^{2}$ region compared to the PPD model, and they fail to satisfy the generalized GT relation Shintani:2018ozy ; Tsuji:2023llh .
In this work, we propose a simple method to remove the $\pi N$-state contamination from the $F_{P}$ and $G_{P}$ form factors given by the standard ratio method in lattice QCD Sasaki:2025qro . The application of the novel method will enable the precise determination of the induced pseudoscalar coupling ($g_{P}^{\ast}$) and the pion-nucleon coupling ($g_{\pi NN}$) from existing data calculated at two lattice spacings, since these quantities are defined using the $F_{P}$ form factor Thomas:2001kw as
@@
$g_{\pi NN}$ $=\lim_{q^{2}\rightarrow\infty}(q^{2}+M_{\pi}^{2})\frac{F_{P}(q^{2})}{2F_{\pi}},$
@@
This paper is organized as follows: in Sec. II, after a brief introduction to the standard ratio method for calculating the nucleon form factors in lattice QCD simulations, we describe our proposal to remove $\pi N$-state contamination from two specific nucleon form factors $F_{P}(q^{2})$ and $G_{P}(q^{2})$ (hereafter referred to as the leading $\pi N$ subtraction method). Section III presents the numerical results of the $F_{P}$ and $G_{P}$ form factors obtained by the leading $\pi N$ subtraction method with the datasets generated in Refs. Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut . We then determine the values of $g_{P}^{\ast}$ and $g_{\pi NN}$. Finally, we close with a summary in Sec. IV.
In this paper, the nucleon matrix elements are given in the Euclidean metric convention. $\gamma_{5}$ is defined by $\gamma_{5}\equiv\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}=-\gamma_{5}^{M}$, which has the opposite sign relative to that in the Minkowski convention ($\vec{\gamma}^{M}=i\vec{\gamma}$ and $\gamma^{M}_{0}=\gamma_{4}$) adopted in the Particle Data Group. The sign of all the form factors is chosen to be positive. The Euclidean four-momentum squared $q^{2}$ corresponds to the spacelike momentum squared as $q_{M}^{2}=-q^{2}<0$ in Minkowski space.
Table 1: Details of gauge ensembles: the $\beta$ value, lattice volume ($L^{3}\times T$), spatial extent (LaLa), the simulated pion mass ($M_{\pi}$), the number of configurations ($N_{\mathrm{conf}}$) and the total number of the measurements ($N_{\mathrm{meas}})$ using on the all-mode-averaging technique Blum:2012uh ; Shintani:2014vja ; vonHippel:2016wid to be analyzed for studying the isovector nucleon form factors in our previous works Shintani:2018ozy ; Tsuji:2023llh ; Ishikawa:2021eut ; Tsuji:2022ric and also this study.
$\beta$ Label $L^{3}\times T$ LaLa [fm] $M_{\pi}$ [GeV] $N_{\mathrm{conf}}$ $N_{\mathrm{meas}}$
1.82 PACS10/L128 $128^{3}\times 128$ 10.8 0.135 20 $O(2500)$-$O(10000)$
PACS5/L64 $64^{3}\times 64$ 5.9 0.138 100 $O(25000)$-$O(200000)$
2.00 PACS10/L160 $160^{3}\times 160$ 10.1 0.138 19 $O(5000)$-$O(58000)$
@@ II.1 Standard ratio method
where the local current $\widetilde{J}$  [^2] Hereafter, the currents and the form factors with and without tildes indicate bare and renormalized ones. located at the time slice $t$ denotes the axial-vector current ($J=A_{\alpha}$) or the pseudoscalar current ($J=P$), carrying the momentum transfer $\bm{q}=\bm{p}-\bm{p}^{\prime}$. The projection operator ${\cal P}_{+}=\frac{1+\gamma_{4}}{2}$ appearing in Eq. (7) can eliminate contributions from the opposite-parity state for the case of $|\bm{p}|=0$ Sasaki:2001nf ; Sasaki:2005ug , while the projection operator ${\cal P}^{5z}={\cal P}_{+}\gamma_{5}\gamma_{3}$ appearing in Eq. (8) means that the $z$ direction is chosen as the polarization direction.
In a conventional way to extract the nucleon form factors, the following ratios are constructed by appropriate combinations of two-point and three-point functions with the source-sink separation ($t_{\mathrm{sep}}\equiv t_{\mathrm{sink}}-t_{\mathrm{src}}$Hagler:2003jd ; Gockeler:2003ay as
@@
Figure 1: Using the standard ratio method, the values of $F_{A}$ (upper-left panel), $2M_{N}F_{P}$ (upper-right panel) and $\widetilde{G}_{P}$ (lower panel) calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven nonzero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$.
The ratio $\mathcal{R}_{J}^{5z}(t;\bm{q})$ gives the following asymptotic values including the respective bare form factors in the asymptotic region Sasaki:2007gw , as follows:
@@
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh the correlation functions involving the $A_{4}$ current are not taken into account for the calculation of the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ form factors. Instead of using Eq. (13), kinematically different types of the three-point functions (12) are used to extract the two independent form factors $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$, since the $z$ direction is chosen as the polarization direction by the projection operator ${\cal P}^{5z}$ Sasaki:2007gw . Indeed, the longitudinal momentum ($q_{3}$) dependence explicitly appears in Eq. (12). Therefore, in the transverse case ($i=1$ or 2), ${\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ only contains the $\widetilde{F}_{P}(q^{2})$ form factor, while ${\cal R}^{5z}_{A_{3}}(t,{\bm{q}})$ contains both $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ contributions as shown in Eq. (12). Even in the longitudinal case ($i=3$), the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ can be separately determined due to the fact that there are two types of kinematics: $q_{3}=0$ or $q_{3}\neq 0$. For the former case, ${\cal R}^{5z}_{A_{3}}(t,{\bm{q}})$ only contains the $\widetilde{F}_{A}(q^{2})$. Therefore, the subtracted correlator ratio $\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ that only includes the $\widetilde{F}_{P}(q^{2})$ contribution can be defined as
$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ $\equiv{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})-\delta_{i3}{\cal R}^{5z}_{A_{3}}(t,{\bm{q}}_{0})$ (15)
@@
Finally, we recall that the quark local currents on the lattice receive finite renormalizations relative to their continuum counterparts in general. The renormalized values of the form factors require the renormalization factors $Z_{O}\ (O=A,P)$, as follows:
@@
where the renormalization factors are defined through the renormalization of the quark local currents, $A_{\alpha}=Z_{A}\widetilde{A}_{\alpha}$ and $P=Z_{P}\widetilde{P}$. The renormalization factor $Z_{A}$ is scale independent, while $Z_{P}$ depends on the renormalization scale. To make a comparison with the experimental values, two form factors, $F_{A}$ and $F_{P}$, will be properly renormalized with $Z_{A}$, which is determined by the Schrödinger functional method. On the other hand, the pseudoscalar form factor is provided only as a bare quantity as indicated by $\widetilde{G}_{P}$ in this study.
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , we adopted the strategy that the smearing parameters of the nucleon interpolating operator are highly optimized to eliminate as much as possible the contribution of excited states in the nucleon two-point function. Figure 1 shows the values of $F_{A}$ (upper-left panel), $2M_{N}F_{P}$ (upper-right panel) and $\widetilde{G}_{P}$ (lower panel) computed on the $160^{4}$ lattice (PACS10/L160) with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven nonzero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. Although $F_{A}(q^{2})$ shows a clear plateau and no $t_{\mathrm{sep}}$ dependence, suggesting that the ground-state dominance is manifest thanks to the elaborate tuning of the sink and source smearing functions, the excited-state contributions are not completely eliminated in $F_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$.